OpenStudy (anonymous):

Jessica's bank contains 18 quarters and dimes, of which q are quarters. Find the total value of the coins in dollars.

OpenStudy (anonymous):

@hello1213 Is it 18 Quarters and Dimes combined ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got confused to with the way they worded it but yes combined there are 18 quarters and dimes.

OpenStudy (anonymous):

One dime = $ 0.1 One Quarter = $ 0.25 No. of Quarters = q No. of Dimes = 18-q

OpenStudy (anonymous):

It seems to be One Equation with two variables :S

OpenStudy (anonymous):

so how would i set this up .... would i say q + 18-q= 18

OpenStudy (anonymous):

wait nevermind that would get me to 18=18

OpenStudy (anonymous):

So Cost of q quarters = 0.25q Cost of 18-q dimes = 1.8 - 0.1q Total Cost = 1.8 - 0.1q + 0.25q = 1.8 + 0.15q We have two variables here, q and the total cost.

OpenStudy (anonymous):

ohhh wow thats easier then i thought... thank you!

OpenStudy (anonymous):

:) Your Welcome